OpenStudy (anonymous):

How much will a 6 foot pipe hold if it has a circumference of 37.68 inches?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (ivancsc1996):

I think it has to do with preassure on the pipes.

OpenStudy (campbell_st):

use the circumference information and the formula for circumference of a circle to find the radius. when you have that you can find the volume of the pipe

OpenStudy (anonymous):

I wasn't asking for the answer(: V= 3.14 * r ^2 * h And I know C= 2 * 3.14 * r But I don't know how to go about this

OpenStudy (anonymous):

Do I use the 6 foot height and solve from there, and the circumference and solve from there, and use the information together?

OpenStudy (campbell_st):

and you also know C = 37.68

OpenStudy (campbell_st):

the process is find the radius 1st... before doing anything with volume

OpenStudy (campbell_st):

well you know \[37.68 = 2 \times \pi \times r\] solve for r.

OpenStudy (skullpatrol):